THE FORMATION OF SH^{+} VIA S^{2+} + H_{2} IN X-RAY DOMINATED REGIONS

N. P. ABEL, Department of Physics, University of Cincinnati, Cincinnati, OH; S. R. FEDERMAN, Department of Physics & Astronomy, University of Toledo, Toledo, OH; and P. C. STANCIL, Department of Physics & Astronomy and the Center for Simulation Physics, University of Georgia.

A. Dalgarno first suggested thirty-two years ago that doubly-ionized X^{2+} elements reacting with H_{2} could lead to the formation of molecules in regions with a significant X-ray flux. However, few investigations of the effects of X^{2+} on the formation of molecules have actually been conducted. Recent studies have found the reaction rate coefficients to be orders of magnitude higher than previously thought. This fact, combined with a renewed interest in the physics and chemistry of X-ray Dominated Regions (XDRs), points to a need to study the effects of X^{2+} on molecule formation. We investigated how S^{2+} + H_{2} affects the formation of SH^{+} for atomic gas exposed to an X-ray continuum typical of an Active Galactic Nucleus (AGN). As long as the fraction of S^{2+} + H_{2} reactions which lead to SH^{+} formation exceeds a few percent, this process will be the dominant formation pathway for SH^{+}. This potentially has major consequences for our understanding of sulfur chemistry in galaxies where the energetics is dominated by AGN activity.